# Phase-Space Engineering and Dynamical Long-Range Order in Memcomputing

### v1

@@ Abstract
Digital Memcomputing machines (DMMs) are dynamical systems with memory (time non-locality) that have been designed to solve combinatorial optimization problems. Their corresponding ordinary differential equations depend on a few hyper-parameters that define both the system’s relevant time scales and its phase-space geometry. Using numerical simulations on a prototypical DMM, we analyze the role of these physical parameters in engineering the phase space to either help or hinder the solution search by DMMs. We find that the DMM explores its phase space efficiently for a wide range of parameters, aided by the long-range correlations in their fast degrees of freedom that emerge dynamically due to coupling with the (slow) memory degrees of freedom. In this regime, the time it takes for the system to find a solution scales well as the number of variables increases. When these hyper-parameters are chosen poorly, the system navigates its phase space far less efficiently. However, we find that, in many cases, dynamical long-range order (DLRO) persists even when the phase-space exploration process is inefficient. DLRO only disappears if the memories are made to evolve as quickly as the fast degrees of freedom. This study points to the important role of memory and hyper-parameters in engineering the DMMs’ phase space for optimal computational efficiency.
@@ I Introduction
Memcomputing represents a new type of computing paradigm where memory (i.e., time non-locality) is employed as a computing tool [1, 2, 3]. In particular, digital memcomputing machines (DMMs) [4, 5], which map a finite string of symbols into a finite string of symbols, are physically represented by dynamical systems that leverage memory to solve combinatorial optimization problems. The ensuing nonlinear ordinary differential equations (ODEs) couple fast degrees of freedom (DOFs) –representing the variables of the problem to solve– to slow, memory DOFs. In addition, the DMMs’ phase space can be engineered so that the only attractors of the dynamics are the solutions of the problem to solve, and all other critical points are saddle points [3].
Several studies have shown that numerical simulations of DMMs’ equations of motion can efficiently solve a wide range of complex combinatorial optimization problems [4, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] (see also industrial case studies at [18]). However, much less attention has been devoted to better engineering the phase space of DMMs by varying some of the hyper-parameters that enter the ODEs of DMMs (see also Sec. III). In fact, we expect these parameters to change the curvatures of both the stable and unstable manifolds of all the critical points in the phase space, thus favoring or hindering the solution search of DMMs.
@@
1. There exists a wide range of viable DMM hyper-parameters where the solution search is efficient. This phase of solvability is related to the existence of a full phase of dynamical long-range order (DLRO) in the fast DOFs that is induced by memory. By “long-range” we mean that the correlations between the fast DOFs may involve a large number of these DOFs, possibly their entire set  [^1] In the presence of an underlying spatial structure, e.g., a spin glass on a lattice, DLRO is directly related to spatial long-range correlations [29]. However, in many combinatorial optimization problems, such a structure is missing. Nonetheless, the concept of DLRO is still well-defined if we consider the size of fast DOF avalanches generated during dynamics. Namely, we consider systems exhibiting collective motion of a large number of fast DOFs to have DLRO [46].. It also indicates that DMMs are robust against small perturbations of the hyper-parameters within this phase.
2. Despite there being a wide range of hyper-parameters that yield DLRO, only a subset of these parameters is accompanied by scale-invariance. While DLRO is certainly required to navigate highly non-convex landscapes efficiently, scale-invariance does not seem to be always necessary.
3. When the memory DOFs vary over time scales comparable to those of the fast DOFs, the DMM’s ability to solve the problem decreases dramatically. This loss of efficiency originates from the breakdown of DLRO in the fast DOFs, as timescale separation is required to generate long-range correlations dynamically.
@@
The paper is organized as follows. In Sec. II, we analyze the role that memory plays in DMMs, both establishing DLRO and shaping/navigating the DMM phase space. In Sec. III, we introduce the 3-SAT problem and the full DMM equations which we simulate to solve hard 3-SAT instances. Finally, in Sec. IV, we perform some numerical analyses to substantiate the claims made above.
@@ II Memory’s Role in Memcomputing
II.1 Dynamical Long-Range Order via Memory
We begin with a general analysis of some generic DMMs’ equations of motion that does not depend on their precise functional form. In fact, apart from the necessity of having two types of DOFs, slow (memory) and fast (variables of the problem to solve), the ODEs of DMMs may be different even for a specified computational problem [3]. That said, we will now provide analytical arguments on how, within a certain region in hyper-parameter space, DLRO in the fast DOFs emerges due to the presence of memory, which aids the system in the solution search process.
@@ II.1 Dynamical Long-Range Order via Memory
where $n=1,2,...,N$ and $m=1,2,...,\tilde{M}$, with $N$ the number of logical variables and $\tilde{M}$ the number of memory variables. Both $C_{m}(\vec{v})\in[0,1]$ and $\gamma\in[0,1]$.
Note that all variables have been continuously relaxed; e.g., $v_{i}\in\{-1,1\}\rightarrow[-1,1]$ (where we take $-1$ and $+1$ to be the booleans FALSE and TRUE in the original, logical problem)  [^3] For the relaxed variables, it is enough to assign $v_{i}>0$ to the logical TRUE, and $v_{i}<0$ to the logical FALSE.. This relaxation is crucial in order to transform the discrete configuration space of the original, logical problem into a continuous phase space on which our DMM can evolve. In Sec. III we will provide explicit functional forms for the functions $A_{n,m}(\vec{v})$, $B_{n,m}(\vec{v})$, and $C_{m}(\vec{v})$. Here, we only assume them to be continuous real-valued functions on the space of $v_{i}$’s.
We will now show how the phenomenon of DLRO in the $v_{i}$’s can emerge in such a system due to the memory variables $x_{m}$’s. Similar calculations have been performed recently in the context of neuromorphic devices [23], spin systems [24], and the biological brain [25]. This process requires that the memory variables evolve more slowly than the logical variables so that the $x_{m}$’s preserve some “memory” of the system’s recent history. In fact, the presence of such memory is necessary in order for DMMs to function properly.
To infer the existence of long-range correlations, we iteratively integrate our dynamics, similar to how a computer would simulate such dynamics numerically. A similar technique has been used in previous works in other contexts, including transport theory [26] and neural dynamics [27]. We find this approach to be much more effective (and feasible) than, e.g., multiple-scale analysis [28] or any other existing semi-analytical techniques. During the iterative integration, we fix the integration timestep to be $\tau_{x}$, the characteristic timescale of the memory variable dynamics, which satisfies $\tau_{x}>\tau_{v}$, with $\tau_{v}$ as the characteristic timescale of the logical variable dynamics.
@@
[…]
@@
Note that this analysis requires that $\beta$ not be too large. If instead $\beta\gtrsim|\sum_{m}\overline{C_{m}^{(0)}(\vec{v})}\,\overline{A_{n,m}^{(1)}(\vec{v})}|$, the effect becomes perturbative (so that many $v_{i}$’s do not correlate strongly nor evolve collectively) precisely when $x_{m}$ evolves too quickly (so it is no longer an effective source of memory on the system’s recent history). Of course, the exact value of $\beta$ which qualifies as “too large” will depend on the functional form of $A_{n,m}(\vec{v})$ and any physical parameters therein (cf. Sec. III).
@@ II.2 Effective Phase Space Navigation via Memory
The long-range correlations we have just described are precisely what enables DMMs to solve combinatorial optimization problems efficiently. These problems are typically plagued by highly non-convex landscapes in their state spaces, and traditional algorithms (which rely only on “local” information in the solution search process) frequently get stuck in local minima. For example, a traditional algorithm that quickly finds a reasonably good configuration (i.e., a low-energy configuration that is not the ground state) may still need to change the value of many (a macroscopic number of) variables to reach the true solution (the ground state). It can then be quite challenging to find a better configuration by updating variables one or a few at a time. Instead, by introducing memory into the system, the (continuously relaxed) logical variables in DMMs become strongly coupled during the solution search so that they can evolve collectively (leveraging the global structure of the problem) towards a logical solution.
@@
Additional insight into the collective dynamics of DMMs has been obtained by means of supersymmetric topological field theory [29, 30], showing that the only low-energy/long-wavelength trajectories in the DMMs’ phase space are a collection of successive instantons. Practically speaking, these instantons correspond to sudden “jumps” in phase space in which many logical variables (even as many as the size of the problem) change suddenly. Instantons are then the underlying expression of DLRO. Importantly, these jumps are between saddle points of decreasing index (with the index being the number of unstable directions of the critical point), until the system eventually converges to a solution (with index 00). During each jump, the system’s state is constrained to lie on an instanton manifold. The union of these manifolds and the critical points between them (collectively called the “composite instanton”) is a proper subset of the entire phase space (in fact it is a CW complex [3]). Therefore, during this characteristic DMM search process, the machine breaks ergodicity. Finally, the dynamics will only stop once a solution (if it exists) is reached (which is a fixed point of the flow field). This behavior has been explicitly shown in several previous studies [29, 31, 3].
However, as we have just discussed, this characteristic DMM search process relies upon the proper choice of hyper-parameters. As we will demonstrate shortly, poorly chosen parameters can dramatically change the effectiveness of a DMM at finding a solution to a given problem. This can occur due to a breakdown of DLRO and/or the influence of strong numerical noise  [^4] We expect a similar limitation in the presence of strong physical noise..
@@ III.2 DMM’s 3-SAT Equations
We now introduce one possible set of DMM equations that we will simulate numerically, which have proven effective (in the correct parameter regime) in solving the above 3-SAT instances [9]  [^5] Note that the equations used in the present work are slightly different than those used in Ref. [9], in the representation of the short- and long-term memory variables (see the SM for more details). These changes do not affect the solution search, further reinforcing that DMMs are quite robust even with respect to structural changes of their flow vector field [3]. ($v_{n},v_{j},$ and $v_{k}$ are the three variables of a single 3-SAT OR gate):
@@
$\dot{x}_{m}^{l}$ =$=\begin{cases}\alpha(C_{m}(v_{i},v_{j},v_{k})-\delta),&C_{m}(v_{i},v_{j},v_{k})\geq\delta,\\ \alpha(C_{m}(v_{i},v_{j},v_{k})-\delta)(x_{l,m}-1),&C_{m}(v_{i},v_{j},v_{k})<\delta,\end{cases}$ (13)
@@
$R_{n,m}(v_{n},v_{j},v_{k})$ =$=\begin{cases}\frac{1}{2}q_{n,m}(1-q_{n,m}v_{n}),&C_{m}(v_{n},v_{j},v_{k})=\frac{1}{2}(1-q_{n,m}v_{n}),\\ 0,&\text{else},\end{cases}$ (15)
$C_{m}(v_{i},v_{j},v_{k})$ $=\frac{1}{2}\min[(1-q_{i,m}v_{i}),(1-q_{j,m}v_{j}),(1-q_{k,m}v_{k})],$ (16)
where $\alpha$, $\beta$, $\gamma$, $\delta$, and $\zeta$ are the hyper-parameters we will analyze in detail in Sec. IV. Note that the 3-SAT instance is solved if $C_{m}<1/2$ for all clauses [9]. When simulating these equations, each instance is initialized with random $v_{i}=\pm 1$ and fixed $x_{m}^{s}=1/2$ and $x_{m}^{l}=1$. These equations correspond to Eq. (1) and Eq. (2), with $A_{n,m}$, $B_{n,m}$, and $C_{m}$ defined as follows:
@@
Notice that in Eqns. (14), (15), and (16), $q_{i,m}$ indicates whether or not a logical variable is negated in the literal appearing in the relevant 3-SAT formula: $q_{i,m}=1$ if $l_{i,m}=v_{i}$ (in clause $m$), and $q_{i,m}=-1$ if $l_{i,m}=\lnot v_{i}$. Since our solution search process is dynamical rather than algorithmic, to make a particular literal $l_{i,m}$ TRUE in the original discrete problem, we must “push” $v_{i}$ towards $q_{i,m}$. Thus, we can also reformulate any literal $l_{i,m}$ as a continuous dynamical variable $q_{i,m}v_{i,m}$. See the SM for a more detailed description of these dynamics.
We emphasize that this is just one possible form of DMM’s dynamics—the specific functional form is not unique, as long as there is timescale separation between the logical DOFs and at least one type of memory DOF to enable DLRO. In fact, different ODE formulations may be more effective for different types of problems  [^6] Note also that in order to accelerate numerical simulations, the parameters in Eqns. (14), (15), and (16) have been chosen so that the dynamics approach critical points in the phase space but never “fall into” one exactly (except for the equilibrium points). This speeds up the approach to equilibrium without qualitatively changing the phase space exploration..
@@
It must be emphasized that there is a crucial difference between the physical parameters in Eqns. (11), (12), and (13) and the numerical parameter $dt_{0}$. Although some $dt_{0}$ tuning is required to control numerical errors, only $\alpha$, $\beta$, $\gamma$, $\delta$, and $\zeta$ influence physical DMM properties such as the geometry of the phase space and the correlations between logical variables. For example, if a DMM was constructed in hardware [40, 16, 17] that evolved according to Eqns. (11) through (16), $dt_{0}$ would not play any role at all. Therefore, since we are most interested in the physical properties of DMMs in this work, we will concentrate only on the physical DMM parameters for the remainder of this manuscript. An additional manuscript is planned for the near future which will address the impact of numerical parameters on DMMs’ problem-solving efficiency.

### v2

# Phase-Space Engineering and Collective Dynamics in Memcomputing

@@ Abstract
Digital Memcomputing machines (DMMs) are dynamical systems with memory (time non-locality) that have been designed to solve combinatorial optimization problems. Their corresponding ordinary differential equations depend on a few hyper-parameters that define both the system’s relevant time scales and its phase-space geometry. Using numerical simulations on a prototypical DMM, we analyze the role of these physical parameters in engineering the phase space to either help or hinder the solution search by DMMs. We find that the DMM explores its phase space efficiently for a wide range of parameters, aided by the system-wide correlations in their fast degrees of freedom that emerge dynamically due to coupling with the (slow) memory degrees of freedom. In this regime, the time it takes for the system to find a solution scales well as the number of variables increases. When these hyper-parameters are chosen poorly, the system navigates its phase space far less efficiently. However, we find that, in many cases, collective behavior persists even when the phase-space exploration process is inefficient. This behavior only disappears if the memories are made to evolve as quickly as the fast degrees of freedom. This study points to the important role of memory and hyper-parameters in engineering the DMMs’ phase space for optimal computational efficiency.
@@ I Introduction
Memcomputing represents a new type of computing paradigm where memory is employed as a computing tool [1, 2, 3]. Here, “memory” refers to the general time non-local nature of the corresponding dynamical system (rather than the more restrictive meaning of “data storage”). This memory plays a crucial role in enabling memcomputing machines to solve challenging computational problems efficiently. These machines can be fabricated using a variety of materials and devices, including electrical devices, optical elements, etc. [3], so long as a time non-local response is present in the dynamical system [4]. Starting from these experimental realizations one can then write down the corresponding nonlinear ordinary differential equations (ODEs) describing their dynamics. In turn, these ODEs can be simulated on traditional computers even without a hardware realization.
Within the general class of memcomputing machines, digital memcomputing machines (DMMs), map a finite string of symbols into a finite string of symbols, enabling them to solve combinatorial optimization problems [5, 6]. Their internal ODEs couple fast degrees of freedom (DOFs) –representing the variables of the problem to solve– to slow, memory DOFs. The memory variables enable the fast DOFs to become strongly coupled during the solution search; the resulting collective behavior leverages the global structure of the problem to find a logical solution efficiently [7]. In addition, the DMMs’ phase space can be engineered so that the only attractors of the dynamics are the solutions of the problem to solve, and all other critical points are saddle points [3]. Altogether, this makes a DMM a powerful tool when tackling difficult combinatorial optimization problems.
@@
1. There exists a wide range of viable DMM hyper-parameters where the solution search is efficient. This phase of solvability is related to the existence of a region in phase space in which the fast DOFs evolve collectively. This behavior is induced by memory. The fact that a phase exists where solutions can be found efficiently indicates that DMMs are robust against small perturbations of the hyper-parameters within this range.
2. Despite there being a wide range of hyper-parameters that yield collective dynamics, only a subset of these parameters is accompanied by scale-invariance. While collective behavior is certainly required to navigate highly non-convex landscapes efficiently, scale-invariance does not seem to be always necessary.
3. When the memory DOFs vary over time scales comparable to those of the fast DOFs, the DMM’s ability to solve the problem decreases dramatically. This loss of efficiency originates from the breakdown of collective behavior in the fast DOFs, as timescale separation is required to generate system-wide correlations dynamically.
@@
The paper is organized as follows. In Sec. II, we analyze the role that memory plays in DMMs, both establishing collective dynamics in the fast DOFs and shaping/navigating the DMM phase space. In Sec. III, we introduce the 3-SAT problem and the full DMM equations which we simulate to solve hard 3-SAT instances. Finally, in Sec. IV, we perform some numerical analyses to substantiate the claims made above.
@@ II Memory’s role in memcomputing
II.1 Collective dynamics via memory
We begin with a general analysis of some generic DMMs’ equations of motion that does not depend on their precise functional form. In fact, apart from the necessity of having two types of DOFs, slow (memory) and fast (variables of the problem to solve), the ODEs of DMMs may be different even for a specified computational problem [3]. That said, we will now provide analytical arguments on how, within a certain region in hyper-parameter space, the fast DOFs behave collectively due to the presence of memory, which aids the system in the solution search process.
@@ II.1 Collective dynamics via memory
where n=1,2,,Nn=1,2,...,N and m=1,2,,Mm=1,2,...,M, with $N$ the number of logical variables and $M$ the number of memory variables. Both $C_{m}(\vec{v})\in[0,1]$ and $\gamma\in[0,1]$.
Note that all variables have been continuously relaxed; e.g., $v_{i}\in\{-1,1\}\rightarrow[-1,1]$ (where we take 1-1 and +1+1 to be the booleans FALSE and TRUE, respectively, in the original, logical problem). For the relaxed variables, it is enough to assign $v_{i}>0$ to the logical TRUE, and $v_{i}<0$ to the logical FALSE. This relaxation is crucial in order to transform the discrete configuration space of the original, logical problem into a continuous phase space on which our DMM can evolve. In Sec. III we will provide explicit functional forms for the functions $A_{n,m}(\vec{v})$, $B_{n,m}(\vec{v})$, and $C_{m}(\vec{v})$. Here, we only assume them to be continuous real-valued functions on the space of $v_{i}$’s.
We will now show how the $v_{i}$’s can evolve collectively in such a system due to the memory variables $x_{m}$’s. Similar calculations have been performed recently in the context of neuromorphic devices [10], spin systems [7], and the biological brain [11]. This process requires that the memory variables evolve more slowly than the logical variables so that the $x_{m}$’s preserve some “memory” of the system’s recent history. In fact, the presence of such memory is necessary in order for DMMs to function properly.
To infer the existence of collective behavior, we iteratively integrate our dynamics, similar to how a computer would simulate such dynamics numerically. A similar technique has been used in previous works in other contexts, including transport theory [12] and neural dynamics [13]. We find this approach to be much more effective (and feasible) than, e.g., multiple-scale analysis [14] or any other existing semi-analytical techniques. During the iterative integration, we fix the integration timestep to be $\tau_{x}$, the characteristic timescale of the memory variable dynamics, which satisfies $\tau_{x}>\tau_{v}$, with $\tau_{v}$ as the characteristic timescale of the logical variable dynamics.
@@
with $F^{(l)}(\vec{v})$ either $A_{n,m}(\vec{v})$ or $B_{n,m}(\vec{v})$, $l=0,1,\dots$, and $r\equiv\lfloor\tau_{x}/\tau_{v}\rfloor$, with $\lfloor\cdot\rfloor$ the floor function of the argument. Essentially, Eq. (5) gives the average of some function $F$ in the time interval $l\tau_{x},(l+1)\tau_{x})$ of width $\tau_{x}$. Notice that $x_{m}$ does not change much during time intervals shorter than $\tau_{x}$, so we can treat it as approximately constant during each integration step. Also, we must have $r\gg 1$ for the system to have memory.
@@
Note that this analysis requires that $\beta$ not be too large. If instead $\beta\gtrsim|\sum_{m}\overline{C_{m}^{(0)}(\vec{v})}\,\overline{A_{n,m}^{(1)}(\vec{v})}|$, the effect becomes perturbative (so that many $v_{i}$’s do not correlate strongly nor evolve collectively) precisely when $x_{m}$ evolves too quickly (so it is no longer an effective source of memory on the system’s recent history). Of course, the exact value of $\beta$ which qualifies as “too large” will depend on the functional form of $A_{n,m}(\vec{v})$ and any physical parameters therein (cf. Sec. III). In the Supplemental Material (SM), we elaborate upon this analysis further with a simple example.
@@ II.2 Effective phase space navigation via memory
The dynamical, collective effect we have just described is precisely what enables DMMs to solve combinatorial optimization problems efficiently. These problems are typically plagued by highly non-convex landscapes in their state spaces, and traditional algorithms (which rely only on “local” information in the solution search process) frequently get stuck in local minima [15]. For example, a traditional algorithm that quickly finds a reasonably good configuration (i.e., a low-energy configuration that is not the ground state) may still need to change the value of many (a macroscopic number of) variables to reach the true solution (the ground state). It can then be quite challenging to find a better configuration by updating variables one or a few at a time. Instead, by introducing memory into the system, the (continuously relaxed) logical variables in DMMs become strongly coupled during the solution search so that they can evolve collectively (leveraging the global structure of the problem) towards a logical solution.
@@
Additional insight into the collective dynamics of DMMs has been obtained by means of supersymmetric topological field theory [16, 17], showing that the only low-energy/long-wavelength trajectories in the DMMs’ phase space are a collection of successive instantons. Practically speaking, these instantons correspond to sudden “jumps” in phase space in which many logical variables (even as many as the size of the problem) change suddenly. Instantons are then the underlying expression of collective behavior in DMMs. Importantly, these jumps are between saddle points of decreasing index (with the index being the number of unstable directions of the critical point), until the system eventually converges to a solution (with index 0). During each jump, the system’s state is constrained to lie on an instanton manifold. The union of these manifolds and the critical points between them (collectively called the “composite instanton”) is a proper subset of the entire phase space (in fact it is a CW complex [3]). Therefore, during this characteristic DMM search process, the machine breaks ergodicity. Finally, the dynamics will only stop once a solution (if it exists) is reached (which is a fixed point of the flow field). This behavior has been explicitly shown in several previous studies [16, 18, 3].
However, as we have just discussed, this characteristic DMM search process relies upon the proper choice of hyper-parameters. As we will demonstrate shortly, poorly chosen parameters can dramatically change the effectiveness of a DMM at finding a solution to a given problem. This can occur due to a breakdown of collective behavior in the fast DOFs and/or the influence of strong numerical noise [19]. In fact, a similar limitation in the presence of strong physical noise has been recently shown in Ref. [20].
@@ III.2 DMM’s 3-SAT equations
We now introduce one possible set of DMM equations that we will simulate numerically, which have proven effective (in the correct parameter regime) in solving the above 3-SAT instances [25]  [^2] Note that the equations used in the present work are slightly different from those used in Ref. [25], in the representation of the short- and long-term memory variables (see the SM for more details). These changes do not affect the solution search, further reinforcing that DMMs are quite robust even with respect to structural changes of their flow vector field [3].:
@@
$\dot{x}_{m}^{l}$ $=\begin{cases}\alpha(C_{m}(v_{i},v_{j},v_{k})-\delta),&C_{m}(v_{i},v_{j},v_{k})\geq\delta,\\ \alpha(C_{m}(v_{i},v_{j},v_{k})-\delta)(x_{l,m}-1),&C_{m}(v_{i},v_{j},v_{k})<\delta,\end{cases}$ (13)
@@
$R_{n,m}(v_{n},v_{j},v_{k})$ $=\begin{cases}\frac{1}{2}q_{n,m}(1-q_{n,m}v_{n}),&C_{m}(v_{n},v_{j},v_{k})=\frac{1}{2}(1-q_{n,m}v_{n}),\\ 0,&\text{else},\end{cases}$ (15)
$C_{m}(v_{i},v_{j},v_{k})$ $=\frac{1}{2}\min[(1-q_{i,m}v_{i}),(1-q_{j,m}v_{j}),(1-q_{k,m}v_{k})].$ (16)
$v_{i},v_{j},$ and $v_{k}$ are the three variables of a single 3-SAT OR gate, each of which evolves according to Eq. (11). $x^{s}_{m}$ and $x^{l}_{m}$ are the short- and long-term memory DOFs, which increase or decrease depending on the state of the corresponding clause function (only one of each exists for each clause). While not necessary to the memcomputing paradigm itself, previous studies [25] have found that having two types of memory DOFs with the above functional forms aids in the solution search. Finally, $\alpha$, $\beta$, $\gamma$, $\delta$, and $\zeta$ are the physical hyper-parameters. $\beta$ and $\alpha$ influence the rate of change of the short- and long-term memories, respectively (with $\alpha<\beta$). $\gamma$ and $\delta$ are thresholding parameters that must both be less than $1/2$. $\zeta$ influences the degree of “rigidity” in the system, or more precisely, the long-term memories’ relative influence on the system’s rigidity (see Ref. [25] for more discussion on the role of these quantities).
Note that the 3-SAT instance is solved if $C_{m}<1/2$ for all clauses [25]. Once this occurs, the simulation is stopped, and the sign of the logical variables can be read as a boolean string that corresponds to the problem’s logical solution. When simulating these equations, each instance is initialized with random $v_{i}=\pm 1$ and fixed $x_{m}^{s}=1/2$ and $x_{m}^{l}=1$. These equations correspond to Eq. (1) and Eq. (2), with $A_{n,m}$, $B_{n,m}$, and $C_{m}$ defined as follows:
@@
Notice that in Eqns. (14), (15), and (16), $q_{i,m}$ indicates whether or not a logical variable is negated in the literal $l_{i,m}$ appearing in the relevant 3-SAT formula: $q_{i,m}=1$ if $l_{i,m}=v_{i}$ (in clause $m$), and $q_{i,m}=-1$ if $l_{i,m}=\lnot v_{i}$. Since our solution search process is dynamical rather than algorithmic, to make a particular literal $l_{i,m}$ TRUE in the original discrete problem, we must “push” $v_{i}$ towards $q_{i,m}$. Thus, we can also reformulate any literal $l_{i,m}$ as a continuous dynamical variable $q_{i,m}v_{i,m}$. See the SM for a more detailed description of these dynamics.
We emphasize that this is just one possible form of DMM’s dynamics—the specific functional form is not unique, as long as there is timescale separation between the logical DOFs and at least one type of memory DOF. In fact, different ODE formulations may be more effective for different types of problems  [^3] Note also that in order to accelerate numerical simulations, the parameters in Eqns. (11), (12), and (13) have been chosen so that the dynamics approach critical points in the phase space but never “fall into” one exactly (except for the equilibrium points). This speeds up the approach to equilibrium without qualitatively changing the phase space exploration..
Figure 1: (a) State space diagram in hyper-parameter space $\{\beta,\zeta\}$, showing the maximum size 3-SAT ($\alpha_{r}=4.3$) instance $N_{max}$ that a DMM can solve within a fixed number of integration steps $T_{median}$ with fixed $dt_{0}=1.0$. The number of logical variables, NN, is increased by $10$ at a time until $T_{median}$ exceeds $10^{6}$ steps. We find a wide region in phase space in which moderately large $N$ instances ($\sim 400-600$) can be solved quickly without tuning $dt_{0}$ (which scales the adaptive time step). Colored dots and stars correspond to parameter choices depicted in (b). (b) Scalability plots of the median number of steps to find a solution $T_{median}$ as a function of number of logical variables NN. For two sets of parameters ($\{\beta,\zeta\}=\{\beta_{opt},\zeta_{opt}\}=\{79,0.063\}$ and $\{6.0\beta_{opt},1.5\zeta_{opt}\}$), the $T_{median}(N)$’s are fitted well by polynomials ($\sim N^{1.65}$ and $\sim N^{3.18}$, respectively). Outside the phase of “good” parameters, $T_{median}(N)$’s are fitted well by exponentials ($\sim e^{0.12N}$, $\sim e^{0.07N}$, and $\sim e^{0.04N}$ for $\{\beta,\zeta\}=\{\beta_{opt},\zeta_{opt}/10\}$, $\{\beta_{opt}/10,\zeta_{opt}\}$, and $\{10\beta_{opt},\zeta_{opt}\}$, respectively). Here, $dt_{0}$ is tuned at each data point to control numerical errors; everywhere, it takes a value between $0.01$ and $1.0$. In comparison with the memcomputing approach, the inset shows that $T_{median}$ scales exponentially for three different physics-inspired solvers, including coherent Ising machines [28], gain-dissipative dynamics [29], and simulated bifurcation machines [30]. In both figures, batches of $100$ instances are run in parallel, and the batch terminates when at least half of all instances are solved. $\{\alpha/\beta,\beta,\gamma,\delta,\zeta\}_{opt}=\{0.45,79,0.36,0.080,0.063\}$ in both (a) and (b).
@@
It must be emphasized that there is a crucial difference between the physical parameters in Eqns. (11), (12), and (13) and the numerical parameter $dt_{0}$. Although some $dt_{0}$ tuning is required to control numerical errors, only $\alpha$, $\beta$, $\gamma$, $\delta$, and $\zeta$ influence physical DMM properties such as the geometry of the phase space and the correlations between logical variables. For example, if a DMM was constructed in hardware [32, 33, 34] that evolved according to Eqns. (11) through (16), $dt_{0}$ would not play any role at all, as there is no longer any notion of an “integrated trajectory” which approximates the true one. In all other ways, physical noise is expected to influence the dynamical behavior of a memcomputing machine built in hardware in the same ways as numerical noise [20]. A key example in which either numerical or physical noise could play the same role will be discussed in Sec. IV.4.
Since we are most interested in the physical properties of DMMs in this work, we will concentrate only on the physical DMM parameters for the remainder of this manuscript.